# Testing the Shock-cooling Emission Model from Star-disk Collisions for Quasi-periodic Eruptions

### v3

@@ III Constraints for the Model
where $R_{\rm{g}}=\frac{GM_{\rm{h}}}{c^{2}}$ is the gravitational radius of the BH and $c$ is the speed of light in vacuum. In order to simplify the derivation, we define the following notation: $r_{\star}=R_{\star}/\rm{R}_{\rm{\odot}}$, $M_{\rm{6}}=M_{\rm{h}}/10^{6}\ \rm{M}_{\rm{\odot}}$, $P_{\rm{1}}=P_{\rm{orb}}/10\ \rm{hr}$ and $\alpha_{-1}=\alpha/0.1$, where $\rm{R}_{\rm{\odot}}$ and $\rm{M}_{\rm{\odot}}$ are the solar radius and mass, $\alpha$ is the dimensionless disk viscosity parameter and $R_{\star}$ is the radius of the star.
@@ III.1 The Disk
For the inner region of the disk where it is radiation-pressure and radiative-cooling dominated, the scale height of the disk is given by (also see Frank et al., 2002; Linial and Metzger, 2023),
@@
where $v_{\rm{k}}=\sqrt{\frac{GM_{\rm{h}}}{R}}$ is the Kepler velocity at radius $R$, $\kappa$ is the opacity, $\dot{M}$ is the mass accretion rate at the collision point. Hereafter, we consider that the opacity is dominated by Thompson scattering, thus $\kappa=0.34\ \rm{cm}^{2}\rm{g}^{-1}$. The last step of the above has used $R\gg 2R_{\rm{g}}$. We use the lowercase letter $\dot{m}\equiv\dot{M}/\left(L_{\rm{Edd}}/c^{2}\right)$ to denote the dimensionless accretion rate, where $L_{\rm{Edd}}$ is the Eddington luminosity. Then, the luminosity created by accretion is $L=\eta\dot{m}L_{\rm{Edd}}$, where $\eta$ is the radiation efficiency.
@@
$\Sigma_{\rm{d}}=\frac{\dot{M}}{3\pi\nu}=2\times 10^{4}\frac{P_{\rm{1}}}{\alpha_{-1}\dot{m}M_{6}}\ \rm{g/cm^{2}},$ (3)
@@ III.2 Hydrodynamics in Collision
Taking into account $\gamma=\frac{4}{3}$ and the relation $P=\left(\gamma-1\right)\epsilon$, the speed of sound of the immediately post-shock material is
@@ III.3 Duration of Emission
is the surface density of the fireball, $R_{\rm{diff}}=v_{\rm{ej}}t_{\rm{e}}$ is the radius of the fireball at its peak and $t_{\rm{e}}$ is the duration of the eruption. Therefore, one has (also see Linial and Metzger, 2023)
@@ III.4 Luminosity
Next, we consider the energy evolution of the shocked matter. The following Eq. (5), the total internal energy of the immediately post-shock matter, which is dominated by radiation, is $E_{\rm{i,0}}=M_{\rm{sh}}v_{\rm{k}}^{2}$. Note that it is on the same order of magnitude as the total kinetic energy of the shocked matter $E_{\rm{k}}=\frac{1}{2}M_{\rm{sh}}v_{\rm{k}}^{2}$ (Linial and Metzger, 2023).
@@
$=3.9\times 10^{44}r_{\star}^{\frac{5}{3}}\dot{m}^{-\frac{1}{6}}\alpha_{\rm{-1}}^{-\frac{1}{2}}M_{\rm{6}}^{\frac{1}{3}}\ \rm{erg/s}.$ (9)
@@
$L_{\rm{p}}=\frac{E_{i}\left(R_{\rm{diff}}\right)}{t_{\rm{e}}}=6.1\times 10^{41}M_{\rm{6}}\left(\frac{r_{\star}}{P_{\rm{1}}}\right)^{\frac{2}{3}}\dot{m}^{\frac{1}{3}}\ \rm{erg/s}.$ (10)
@@ III.6 Infer $r_{\star}$ from Observables
Eqs. (8), (10) and (11) give the three observables $(t_{\rm{PQE}}$, $L_{\rm{p}}$ and $L_{\rm{Q}})$ from the shock-cooling emission model induced by star-disk collisions. Here, we attempt to investigate what $r_{\star}$ is allowed or required by the model to satisfy these observables.
Notice that $\alpha$ and $\eta$ are the two fudge “factors” that describe the working mechanism of the accretion disk, and $\dot{m}$ is one of the disk properties; all others are observable, or measurable in principle. One can get rid of $\dot{m}$ and express $r_{\star}$ in terms of observables in two ways. First, one can infer $\dot{m}$ from Eq. (10), and plug it into Eq. (8) to get
@@
where $L_{\rm{Q,41}}=L_{\rm{Q}}/\left(10^{41}\ \rm{erg/s}\right)$ and $L_{\rm{p,42}}=L_{\rm{p}}/\left(10^{42}\right$. $\left.\rm{erg/s}\right)$. In this way, we obtain two constraints on $r_{\star}$, both of which must be satisfied, and each contains one fudge factor only.
@@ III.7 Radiation Temperature
When the star-disk collisions occur, the material shocked by the collision will be radiation-dominated and optically thick. Firstly, the radiation energy density is expressed as $u_{\gamma}=L_{\rm{p}}/[4\pi R_{\rm{diff}}^{2}v_{\rm{ej}}]$ (Linial and Metzger, 2023). Secondly, assuming that the radiation of QPEs is dominated by blackbody radiation, $u_{\gamma}=aT_{\rm{p}}^{4}$, where $a$ is the radiation constant, then $T_{\rm{p}}$ can be expressed in terms of various observables as (also see Linial and Metzger, 2023)
@@ III.8 Constraint from Stellar Tidal Disruption
Now we consider another constraint on $r_{\star}$. The stellar orbit must be sufficiently wide; otherwise the star would be tidally disrupted. A critical orbital radius is the so-called tidal disruption radius (e.g., Rees, 1988)
@@
Plugging Eq. (1) in, this can be rearranged to put a constraint on the stellar radius: $r_{\star}<2.4\ n^{-1}P_{\rm{1}}^{\frac{2}{3}}m_{*}^{\frac{1}{3}}$. For a normal star, one can adopt the mass-radius relationship for high-mass zero-age-main-sequence stars : $r_{\star}=m_{*}^{0.6}$ (Kippenhahn et al., 2013). However, as was mentioned in Linial and Metzger (2023); Linial and Sari (2023); Linial and Quataert (2024) and was shown in simulation by Yao et al. (2025), the star may get heated and inflated during collisions. The increased $r_{\star}$ would cause $R_{\rm{t}}$ in Eq. (15) to be larger, making it easier to be disrupted. Then the condition of $R>nR_{\rm{t}}$ would exclude a boarder parameter space.
@@ IV.2 Other QPE Sources
eRO-QPE3 and eRO-QPE4 fare much better in our test, as were shown in the lower panels of Figure 2. The best case is eRO-QPE3, whose allowed parameter space is $r_{\star}\sim 1$ (i.e., a sun-like star) and $M_{\rm{h}}\lesssim 10^{7}\ M_{\rm{\odot}}$, and it is free from any tidal stripping. eRO-QPE4 also has an allowed parameter space, and taking into account the tidal stripping limit, that space shrinks to $r_{\star}\sim 2$ and $M_{\rm{h}}\sim 10^{8}\ M_{\rm{\odot}}$.
@@ IV.4 Comparison of $T_{\rm{p}}$
Figure 4: The comparison of the peak radiation temperatures $T_{\rm{p}}$ predicted by the shock-cooling emission model from star-disk collisions (solid lines) and those observed (crosses with error bars) for the sample. The predicted $T_{\rm{p}}$ is calculated from Eq. (14). The horizontal error bars of the observed $T_{\rm{p}}$ correspond to the estimated $M_{\rm{h}}$ range of each individual source, respectively.
We also compare the predicted peak radiation temperature $T_{\rm{p}}$ (Eq. 14) with the observed ones, which are shown in Figure 4. It is obvious that, except for ASASSN-14ko, the predicted $T_{\rm{p}}$’s for all other sources are one order of magnitude lower than the observed values.
@@ V.1 Orbital Inclination Effect
In this case, the relative velocity is $2v_{\rm{k}}$, the effective length of the stellar trajectory is $\approx 2\pi R$ and the cross section is $\pi R_{\star}\times\min\left[R_{\star},h\right]$. From Eq. (2), one has $h\approx 3\dot{m}M_{\rm{6}}\ R_{\rm{\odot}}$. Considering two regimes $h>R_{\star}$ and $h<R_{\star}$, respectively, one could derive the duration and the peak luminosity, similarly to in $\S$ 3.3 - 3.4, as
@@
Based on Eqs. (17-18) and using Eq.(11), one can derive two solutions of $r_{\star}$ in terms of observables, similarly to Eqs. (12-13). One then checks whether the two solutions have overlapping region, considering a numerical range of $\left(0.01,0.1\right)$ for both $\alpha$ and $\eta$, and whether they satisfy the tidal constraint of Eq. (16), similarly to Figures 1-3.
@@
Therefore, the extremely retrograde orbit is indeed a feasible solution to four sources. However, one should note that for a random orientation of the stellar orbit, the differential probability distribution of $i$ is $\propto\sin i$, such that the probability for $i\in\left(\frac{3\pi}{4},\pi\right)$ is $\approx 15\%$ (also see Suzuguchi and Matsumoto, 2025). Thus, the chances for this most optimistic case are low.
@@ V.2 Stream-disk Collisions
As long as the stream’s column density is larger than the disk surface density: $\Sigma_{\rm{s}}=\frac{\Delta M_{\star}}{\pi R_{\star}\Delta R}\gtrsim\Sigma_{\rm{d}}$ (this requires $P_{\rm{1}}\lesssim 6$, while other parameters are fixed at their fiducial values; see Eq. 18 of Yao et al., 2025) where $\Delta M_{\star}$ is the stream’s mass, the stream-disk collision would shock more disk material and then power a brighter and longer-lived flare than the star-disk collision.
In this scenario, the framework in $\S$ III for constraining the shock-cooling emission still applies. To get the constraints on the stream’s cross section, we can just substitute all term of $R_{\star}^{2}$ in equations there with $R_{\star}\Delta R$, and drop off the tidal disruption constraint Eq. (16). In other words, one can interpret $r_{\star}$ in Eqs. (12-13) as the square root of this cross section in unit of $R_{\odot}$.
After removing the green and black horizontal lines in Figures 1-3, one can see that four sources within our sample (i.e., GSN 069, eRO-QPE2, eRO-QPE3 and eRO-QPE4) have the allowed cross-section size that can simultaneously reproduce the observed $L_{\rm{p}}$, $t_{\rm{e}}$ and $L_{\rm{Q}}$. The rest does not pass our tests on radiative properties. We conclude that, under the condition of $\Sigma_{\rm{s}}\gtrsim\Sigma_{\rm{d}}$, the stream-disk collision can work for the majority of the QPE sources (excluding the two long-period QPE-like events).
On the other hand, if $\Sigma_{\rm{s}}<\Sigma_{\rm{d}}$ (i.e., $P_{\rm{1}}>6$), the reverse shock propagating into the debris stream might be strong enough and the shocked stream may dominate the emission and power the eruption. In this case, the extended radial size of the stream leads to a significant dispersion in its orbital energy. Thus, different segments of the stream arrive the disk plane at markedly different times. The spread of arrival times : $\Delta t_{\rm{s}}=\left[P_{\rm{orb}}\left(R+\Delta R\right)-P_{\rm{orb}}\left(R-\Delta R\right)\right]/2\approx\frac{3}{2}\frac{\Delta R}{R}P_{\rm{orb}}\approx 6.5m_{\star}^{\frac{1}{2}}M_{\rm{6}}^{-\frac{5}{6}}\left(\frac{P_{\rm{1}}}{6}\right)^{\frac{4}{3}}\ \rm{hr}$, would represent the effective collision time scale, which might be comparable or even larger than the expansion time scale (cf. Eq. 8) of the shocked stream material.
@@
V.3 Break-out Emission
Besides the shock-cooling emission model, in analogy with a supernova explosion, the star-disk collision may give rise to a break-out emission (Tagawa and Haiman, 2023). The break-out emission releases the internal energy on a short time scale without the adiabatic energy loss, which is expected to produce a more luminous flare. However, the duration of break-out emission is approximated by its dynamical time scale ($\sim h/v_{\rm{k}}$), which poses a challenge in producing a QPE flare with a typical duration of $1\ \rm{hr}$. Although Tagawa and Haiman (2023) addressed the duration issue by introducing a bow shock profile, more careful checks are required in future work.
@@ V.4 BH as the Collider
The typically observed total radiation energy during one eruption from a QPE source is $E_{\rm{rad}}\simeq 7\times 10^{45}\ \rm{erg}$ (Miniutti et al., 2023b). This would require $m_{*}\sim 10^{3}$, i.e., an intermediate mass BH (IMBH), as the collider. Alternatively, if a QPE is powered by a non-accretion process, such as the shock heating and expansion-cooling of the captured material, the effective radiative efficiency $\eta$ would be lower, thus the required $m_{\star}$ would be even higher.

### v4

@@ III Constraints for the Model
where $R_{\rm{g}}=\frac{GM_{\rm{h}}}{c^{2}}$ is the gravitational radius of the BH and $c$ is the speed of light in vacuum. In order to simplify the derivation, we define the following notation: $r_{\star}=R_{\star}/R_{\rm{\odot}}$, $M_{\rm{6}}=M_{\rm{h}}/10^{6}\ M_{\rm{\odot}}$, $P_{\rm{1}}=P_{\rm{orb}}/10\ \rm{hr}$ and $\alpha_{-1}=\alpha/0.1$, where $R_{\rm{\odot}}$ and $M_{\rm{\odot}}$ are the solar radius and mass, $\alpha$ is the dimensionless disk viscosity parameter, and $R_{\star}$ is the radius of the star.
@@ III.1 The Disk
For the inner region of the disk where it is radiation-pressure and radiative-cooling dominated, the scale height of the disk is given by (also see Frank et al., 2002; Linial and Metzger, 2023)
@@
[…]
@@
$\Sigma_{\rm{d}}=\frac{\dot{M}}{3\pi\nu}=2\times 10^{4}\frac{P_{\rm{1}}}{\alpha_{-1}\dot{m}M_{6}}\ \rm{g\,cm^{\rm{-2}}},$ (3)
@@ III.2 Hydrodynamics in Collision
Taking into account $\gamma=\frac{4}{3}$ and the relation $P=\left(\gamma-1\right)\epsilon$, the speed of sound of the immediately postshock material is
@@ III.3 Duration of Emission
is the surface density of the fireball, $R_{\rm{diff}}=v_{\rm{ej}}t_{\rm{e}}$ is the radius of the fireball at its peak, and $t_{\rm{e}}$ is the duration of the eruption. Therefore, one has (also see Linial and Metzger, 2023)
@@ III.4 Luminosity
Next, we consider the energy evolution of the shocked matter. The following Eq. (5), the total internal energy of the immediately postshock matter, which is dominated by radiation, is $E_{\rm{i,0}}=M_{\rm{sh}}v_{\rm{k}}^{2}$. Note that it is on the same order of magnitude as the total kinetic energy of the shocked matter $E_{\rm{k}}=\frac{1}{2}M_{\rm{sh}}v_{\rm{k}}^{2}$ (Linial and Metzger, 2023).
@@
$=3.9\times 10^{44}r_{\star}^{\frac{5}{3}}\dot{m}^{-\frac{1}{6}}\alpha_{\rm{-1}}^{-\frac{1}{2}}M_{\rm{6}}^{\frac{1}{3}}\ \rm{erg\,s^{\rm{-1}}}.$ (9)
@@
$L_{\rm{p}}=\frac{E_{i}\left(R_{\rm{diff}}\right)}{t_{\rm{e}}}=6.1\times 10^{41}M_{\rm{6}}\left(\frac{r_{\star}}{P_{\rm{1}}}\right)^{\frac{2}{3}}\dot{m}^{\frac{1}{3}}\ \rm{erg\,s^{\rm{-1}}}.$ (10)
@@ III.6 Infer $r_{\star}$ from Observables
Eqs. (8), (10), and (11) give the three observables $(t_{\rm{PQE}}$, $L_{\rm{p}}$ and $L_{\rm{Q}})$ from the shock-cooling emission model induced by star-disk collisions. Here, we attempt to investigate what $r_{\star}$ is allowed or required by the model to satisfy these observables.
Notice that $\alpha$ and $\eta$ are the two fudge “factors” that describe the working mechanism of the accretion disk, and $\dot{m}$ is one of the disk properties; all others are observable or measurable in principle. One can get rid of $\dot{m}$ and express $r_{\star}$ in terms of observables in two ways. First, one can infer $\dot{m}$ from Eq. (10) and plug it into Eq. (8) to get
@@
where $L_{\rm{Q,41}}=L_{\rm{Q}}/\left(10^{41}\ \rm{erg\,s^{\rm{-1}}}\right)$ and $L_{\rm{p,42}}=L_{\rm{p}}/\left(10^{42}\right$. $\left.\rm{erg\,s^{\rm{-1}}}\right)$. In this way, we obtain two constraints on $r_{\star}$, both of which must be satisfied, and each contains one fudge factor only.
@@ III.7 Radiation Temperature
When the star-disk collisions occur, the material shocked by the collision will be radiation dominated and optically thick. Firstly, the radiation energy density is expressed as $u_{\gamma}=L_{\rm{p}}/[4\pi R_{\rm{diff}}^{2}v_{\rm{ej}}]$ (Linial and Metzger, 2023). Secondly, assuming that the radiation of QPEs is dominated by blackbody radiation, $u_{\gamma}=aT_{\rm{p}}^{4}$, where $a$ is the radiation constant, then $T_{\rm{p}}$ can be expressed in terms of various observables as (also see Linial and Metzger, 2023)
@@ III.8 Constraint from Stellar Tidal Disruption
Now we consider another constraint on $r_{\star}$. The stellar orbit must be sufficiently wide; otherwise the star would be tidally disrupted. A critical orbital radius is the so-called tidal disruption radius (e.g., Rees, 1988):
@@
Plugging Eq. (1) in, this can be rearranged to put a constraint on the stellar radius: $r_{\star}<2.4\ n^{-1}P_{\rm{1}}^{\frac{2}{3}}m_{*}^{\frac{1}{3}}$. For a normal star, one can adopt the mass-radius relationship for high-mass zero-age-main-sequence stars : $r_{\star}=m_{*}^{0.6}$ (Kippenhahn et al., 2013). However, as was mentioned in Linial and Metzger (2023), Linial and Sari (2023), and Linial and Quataert (2024) and was shown in simulation by Yao et al. (2025), the star may get heated and inflated during collisions. The increased $r_{\star}$ would cause $R_{\rm{t}}$ in Eq. (15) to be larger, making it easier to be disrupted. Then the condition of $R>nR_{\rm{t}}$ would exclude a boarder parameter space.
@@ IV.2 Other QPE Sources
eRO-QPE3 and eRO-QPE4 fare much better in our test, as shown in the lower panels of Figure 2. The best case is eRO-QPE3, whose allowed parameter space is $r_{\star}\sim 1$ (i.e., a sun-like star) and $M_{\rm{h}}\lesssim 10^{7}\ M_{\rm{\odot}}$, and it is free from any tidal stripping. eRO-QPE4 also has an allowed parameter space, and taking into account the tidal stripping limit, that space shrinks to $r_{\star}\sim 2$ and $M_{\rm{h}}\sim 10^{8}\ M_{\rm{\odot}}$.
@@ IV.4 Comparison of $T_{\rm{p}}$
Figure 4: The comparison of the peak radiation temperatures $T_{\rm{p}}$ predicted by the shock-cooling emission model from star-disk collisions (solid lines) and those observed (crosses with error bars) for the sample. The predicted $T_{\rm{p}}$ is calculated from Eq. (14). The horizontal error bars of the observed $T_{\rm{p}}$ correspond to the estimated $M_{\rm{h}}$ range of each individual source.
We also compare the predicted peak radiation temperature $T_{\rm{p}}$ (Eq. 14) with the observed ones, which are shown in Figure 4. It is obvious that, except for ASASSN-14ko, the predicted $T_{\rm{p}}$’s for all other sources are 1 order of magnitude lower than the observed values.
@@ V.1 Orbital Inclination Effect
In this case, the relative velocity is $2v_{\rm{k}}$, the effective length of the stellar trajectory is $\approx 2\pi R$ and the cross section is $\pi R_{\star}\times\min\left[R_{\star},h\right]$. From Eq. (2), one has $h\approx 3\dot{m}M_{\rm{6}}\ R_{\rm{\odot}}$. Considering two regimes, $h>R_{\star}$ and $h<R_{\star}$, one could derive the duration and the peak luminosity, similarly to in $\S$ 3.3 - 3.4, as
@@
Based on Eqs. (17-18) and using Eq.(11), one can derive two solutions of $r_{\star}$ in terms of observables, similarly to Eqs. (12-13). One then checks whether the two solutions have an overlapping region, considering a numerical range of $\left(0.01,0.1\right)$ for both $\alpha$ and $\eta$, and whether they satisfy the tidal constraint of Eq. (16), similarly to Figures 1-3.
@@
Therefore, the extremely retrograde orbit is indeed a feasible solution to four sources. However, one should note that for a random orientation of the stellar orbit, the differential probability distribution of $i$ is $\propto\sin i$, such that the probability for $i\in\left(\frac{3\pi}{4},\pi\right)$ is $\approx 15\%$ (also see Suzuguchi and Matsumoto, 2026). Thus, the chances for this most optimistic case are low.
@@ V.2 Stream-Disk Collisions
As long as the stream’s column density is larger than the disk surface density, $\Sigma_{\rm{s}}=\frac{\Delta M_{\star}}{\pi R_{\star}\Delta R}\gtrsim\Sigma_{\rm{d}}$ (this requires $P_{\rm{1}}\lesssim 6$, while other parameters are fixed at their fiducial values; see Eq. 18 of Yao et al., 2025) where $\Delta M_{\star}$ is the stream’s mass, the stream-disk collision would shock more disk material and then power a brighter and longer-lived flare than the star-disk collision.
In this scenario, the framework in $\S$ III for constraining the shock-cooling emission still applies. To get the constraints on the stream’s cross section, we can just substitute all term of $R_{\star}^{2}$ in equations there with $R_{\star}\Delta R$ and drop off the tidal disruption constraint Eq. (16). In other words, one can interpret $r_{\star}$ in Eqs. (12-13) as the square root of this cross section in unit of $R_{\odot}$.
After removing the green and black horizontal lines in Figures 1-3, one can see that four sources within our sample (i.e., GSN 069, eRO-QPE2, eRO-QPE3, and eRO-QPE4) have the allowed cross-section size that can simultaneously reproduce the observed $L_{\rm{p}}$, $t_{\rm{e}}$ and $L_{\rm{Q}}$. The rest does not pass our tests on radiative properties. We conclude that, under the condition of $\Sigma_{\rm{s}}\gtrsim\Sigma_{\rm{d}}$, the stream-disk collision can work for the majority of the QPE sources (excluding the two long-period QPE-like events).
On the other hand, if $\Sigma_{\rm{s}}<\Sigma_{\rm{d}}$ (i.e., $P_{\rm{1}}>6$), the reverse shock propagating into the debris stream might be strong enough, and the shocked stream may dominate the emission and power the eruption. In this case, the extended radial size of the stream leads to a significant dispersion in its orbital energy. Thus, different segments of the stream arrive the disk plane at markedly different times. The spread of arrival times, $\Delta t_{\rm{s}}=\left[P_{\rm{orb}}\left(R+\Delta R\right)-P_{\rm{orb}}\left(R-\Delta R\right)\right]/2\approx\frac{3}{2}\frac{\Delta R}{R}P_{\rm{orb}}\approx 6.5m_{\star}^{\frac{1}{2}}M_{\rm{6}}^{-\frac{5}{6}}\left(\frac{P_{\rm{1}}}{6}\right)^{\frac{4}{3}}\ \rm{hr}$, would represent the effective collision time scale, which might be comparable or even larger than the expansion time scale (see Eq. 8) of the shocked stream material.
@@
V.3 Breakout Emission
Besides the shock-cooling emission model, in analogy with a supernova explosion, the star-disk collision may give rise to a breakout emission (Tagawa and Haiman, 2023). The breakout emission releases the internal energy on a short time scale without the adiabatic energy loss, which is expected to produce a more luminous flare. However, the duration of breakout emission is approximated by its dynamical time scale ($\sim h/v_{\rm{k}}$), which poses a challenge in producing a QPE flare with a typical duration of $1\ \rm{hr}$. Although Tagawa and Haiman (2023) addressed the duration issue by introducing a bow shock profile, more careful checks are required in future work.
@@ V.4 BH as the Collider
The typically observed total radiation energy during one eruption from a QPE source is $E_{\rm{rad}}\simeq 7\times 10^{45}\ \rm{erg}$ (Miniutti et al., 2023b). This would require $m_{*}\sim 10^{3}$, i.e., an intermediate-mass BH (IMBH), as the collider. Alternatively, if a QPE is powered by a nonaccretion process, such as the shock heating and expansion cooling of the captured material, the effective radiative efficiency $\eta$ would be lower, thus the required $m_{\star}$ would be even higher.